# Video Prediction Policy: A Generalist Robot Policy with Predictive Visual Representations

### v1

@@ Abstract
Recent advancements in robotics have focused on developing generalist policies capable of performing multiple tasks. Typically, these policies utilize pre-trained vision encoders to capture crucial information from current observations. However, previous vision encoders, which trained on two-image contrastive learning or single-image reconstruction, can not perfectly capture the sequential information essential for embodied tasks. Recently, video diffusion models (VDMs) have demonstrated the capability to accurately predict future image sequences, exhibiting a good understanding of physical dynamics. Motivated by the strong visual prediction capabilities of VDMs, we hypothesize that they inherently possess visual representations that reflect the evolution of the physical world, which we term predictive visual representations. Building on this hypothesis, we propose the Video Prediction Policy (VPP), a generalist robotic policy conditioned on the predictive visual representations from VDMs. To further enhance these representations, we incorporate diverse human or robotic manipulation datasets, employing unified video-generation training objectives. VPP consistently outperforms existing methods across two simulated and two real-world benchmarks. Notably, it achieves a 28.1% relative improvement in the Calvin ABC-D benchmark compared to the previous state-of-the-art and delivers a 28.8% increase in success rates for complex real-world dexterous manipulation tasks.
Figure 1: We observe that the visual representations within the video diffusion model explicitly capture both current and predicted future information. Our Video Prediction Policy, built on these representations, achieves consistent improvements across four benchmarks.
@@ 1 Introduction
Building generalist robot policies capable of solving multiple tasks is an active area of research [8, 36]. Two essential components for constructing such generalist policies are action networks and vision encoders. One line of research focused on developing more advanced action networks, such as employing visual-language pre-trained models [7, 8, 31, 58, 28], training from scratch on diverse robotic datasets [49], incorporating auto-regressive [8] or diffusion architectures [16], and scaling up action networks [33]. Another line of work focuses on learning more effective visual representations [41, 29] for embodied tasks from ego-centric video datasets [21, 20] via contrastive learning [45] or image reconstruction [24].
@@
In this paper, we focus on the visual representation learning. We observe that previous vision encoders, which are pre-trained using contrastive learning between two frames or single-frame reconstruction, fail to adequately capture the physical dynamics inherent in sequential video datasets. Recently, powerful video diffusion models (VDMs) [26, 6, 27, 56, 10], trained with direct video generation objectives on much larger datasets, have demonstrated the ability to generate continuous image sequences and exhibit a strong understanding of the physical world. Inspired by the strong prediction capabilities of VDMs, we hypothesize that they can better capture the physical dynamics within video datasets and inherently contain valuable visual representations that reflect the dynamics and evolution of objects. Moreover, we observe that the visual representations within VDMs are structured with shape $(T,H,W)$, explicitly representing $1$ current step and $(T-1)$ predicted future steps, where $H$ and $W$ correspond to the height and width of single image representation. In contrast, previous vision encoders do not explicitly capture future representations. A comparison is visualized in Figure 2. Based on this distinction, we refer to these latent variables within the video diffusion model as “predictive visual representations”. In the experiment part, we also visualize these predictive representations and find they contain valuable temporal information that reflects the evolution of the physical world.
Our key insight is that these predictive visual representations are highly informative for downstream action learning, as they capture the movement of objects, including the robot itself. Moreover, the ability to predict can be learned from both internet-scale video datasets and various robotic datasets using a consistent video generation loss, enabling us to transfer physical knowledge from large-scale internet datasets to specific robotic systems.
Building on this insight, we introduce the Video Prediction Policy (VPP), which employs a two-stage learning process: First, we finetune a text-guided video prediction (TVP) model [22, 14] from pre-trained video diffusion model [6] using various manipulation datasets, including ego-centric human manipulation [20], open-source robotic datasets [42], and self-collected robot data. This training aims to obtain a controllable video generation model that enhances prediction capabilities in the manipulation domain. Second, we develop a multi-task generalist robot policy conditioned on the predictive representations within the TVP model. Given that the predictive representations in the TVP model remain high-dimensional, with the shape $(T,H,W)$, we employ a video former to distill essential information across spatial and temporal dimensions, followed by a widely used diffusion policy [16] to output actions.
@@
  1. 1.
    To the best of our knowledge, we are the first to leverage the visual representations inside video diffusion models. We find that these representations explicitly express predicted future frames, which we refer to as “predictive visual representations”.
  2. 2.
    We introduce a novel generalist robotic policy, the Video Prediction Policy, by fine-tuning a TVP model in the manipulation domain and then learning actions conditioned on predictive visual presentations in the TVP model.
  3. 3.
    We demonstrate the superior performance of our approach in both simulated and real-world environments, highlighting its versatility.
Figure 2: We use the video diffusion model as a vision encoder to obtain the predictive representations that explicitly express both current and sequential future frames. Previous vision encoders did not have explicit future representations.
@@ 2 Related Works
Visual Representation Learning for Robotics. Self-supervised learning (SSL) techniques, such as contrastive [15, 13], distillation-based  [2, 11], and reconstructive [24, 3], have achieved significant advancements in visual representation learning. Prior research has shown that these SSL techniques enable vision encoders to produce effective representations for embodied AI tasks [55, 54, 43, 46, 12], capturing both high-level semantic and low-level spatial information. Notably, methods like R3M [41], vip [37], VC-1 [38], and Voltron [29] have specifically focused on embodied tasks by innovating pre-training approaches on human manipulation video datasets [20, 21]. However, regardless of the training objective, the learned vision encoders primarily focus on extracting pertinent information from current observations without explicitly predicting future states. In contrast, our Video Prediction Policy leverages predictive representations within video prediction models to explicitly encapsulate both current and predicted future frames.
@@
Future Prediction for Embodied Control Tasks. Existing research also explores the use of future prediction to enhance policy learning [4, 5, 51, 18]. For example, SuSIE [5] conditions its control policy on a predicted future keyframe generated by InstructPix2Pix [9], while UniPi [18] learns the inverse dynamics between two generated frames. These methods typically rely on a single future prediction step to determine actions, which may not accurately capture the complexities of physical dynamics. Additionally, they often operate in raw pixel space, which contains much irrelevant information. GR-1 [51] generates subsequent frames and actions in an autoregressive manner. However, it only generates one image per forward pass, and its prediction quality lags behind that of diffusion-based methods. Furthermore, GR-1 does not leverage pre-trained video foundation models. In contrast, VPP leverages an intermediate representation fine-tuned from a pre-trained video diffusion model, which captures continuous future trajectories to more effectively inform policy learning.
@@
Visual Representation inside Diffusion Models. Diffusion models have achieved remarkable success in the image and video generation tasks [48, 6]. Typically trained as denoisers, diffusion models predict original images from noisy inputs [25]. Research has shown that image diffusion models can also function effectively as vision encoders [53, 34, 23], generating meaningful visual representations. These representations have been proven to be linear-separable for discrimination tasks [53], invaluable for semantic segmentation [34], and versatile for embodied tasks [23]. However, the capabilities of representations within video diffusion models have not been extensively explored. Our findings suggest that variables within VDMs have a unique predictive property not present in other visual representations, making them especially useful for sequential embodied control tasks.
@@ 3 Preliminaries
Diffusion Policy. The diffusion model has also proven effective in action learning, known as diffusion policy [16]. The diffusion policy aims to denoise the action sequence $a_{i}=(\widehat{a}_{i},\widehat{a}_{i+1},...,\widehat{a}_{i+m})$ based on observations $s_{i}$ and instruction. Chi et al. [16] point out that diffusion policy is capable of expressing complex multimodal action distributions and stabilizing training. Recent work [47] further enhances the diffusion policy by incorporating the advanced diffusion transformer (DiT) block [44], a technique we also adopt in the Video Prediction Policy to improve performance.
Figure 3: Video Prediciton Policy first trains a text-guided video prediction (TVP) model for manipulation domain, starting from pre-trained video foundation model. Subsequently, it learns actions based on the predictive representations internal to the TVP model.
@@ 4 Video Prediction Policy
We chose the open-sourced Stable Video Diffusion (SVD) model [6] with 1.5 billion parameters as our foundation. we observe that the open-sourced SVD model conditions only on initial-frame images $s_{0}$ without incorporating language instructions $l$. We augment the model to incorporate CLIP [45] language features $l_{emb}$ using cross-attention layers. Furthermore, we adjust the output video resolution to 16×256×256 to optimize training and inference efficiency. Despite these modifications, we preserve the other components of the original pre-trained SVD framework to retain its core capabilities. We denote this modified version as $V_{\theta}$. In this setup, the initial observation $s_{0}$ is concatenated channel-wise with each predicted frame as a condition. Then model $V_{\theta}$ is trained with diffusion objective, reconstructing the full video sequence $x_{0}=s_{0:T}$ in dataset $D$ from noised samples $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$:
@@
The video prediction objective offers a unified interface that directly generates future visual sequences, enabling the TVP model to harness physical knowledge from diverse datasets. These include internet-based human manipulation datasets $D_{H}$, publicly available robot manipulation data $D_{R}$, and also self-collected datasets $D_{C}$. Given the varying quality and scale of these datasets, we introduce specific coefficients $\lambda$ to appropriately balance the influence of different dataset types:
@@ 4.2 Action Learning Conditioned on Predictive Visual Representation
TVP Model as Vision Encoder. After training the TVP model specifically for manipulation tasks, it can accurately predict future sequences based on image observations and instructions. However, denoising an entire video sequence is highly time-consuming and may lead to open-loop control issues, as discussed in [18]. Moreover, videos in their original pixel format often contain excessive, irrelevant information that can interfere with effective decision-making.
To address these concerns, we employ the video diffusion model primarily as a “vision encoder” rather than a “denoiser” by performing only a single forward step. Our insight is that the first forward step, while not yielding a clear video, still provides a rough trajectory of future states and valuable guidance. This insight is verified in our experiment section and shown in Fig 5. Specifically, we concatenate the current image $s_{0}$ with the final noised latent $q(x_{t^{\prime}}|x_{0})$ (typically white noise) and input this combination into the TVP model. We then directly utilize the latent features $F_{m}\in\mathbb{R}^{T\times W\times H\times C}$ in $m^{th}$ layer of the video diffusion model $V_{\theta}$:
$F_{m}=V_{\theta}(x_{t^{\prime}},l_{emb},s_{0})_{(m)}$ (5)
For a robot with multiple camera views, such as a third-view and a wristed camera, we predict the future for each view independently, denoted as $F^{static}_{m},F^{wrist}_{m}$.
Video Former. These predictive representations within the video diffusion model are still high-dimensional, as they express a sequence of image features. To efficiently aggregate representations across spatial, temporal, and multi-view dimensions, we use a Video Former to consolidate this information into a fixed number of tokens. The Video Former initializes $T\times L$ learnable tokens $Q_{[0:T,0:L]}$, performing spatial-temporal attention on each corresponding frame in the predictive representations, followed by feed-forward layers. Formally, this branch can be expressed as follows where $i$ is the index of frame:
$\begin{split}Q^{\prime}=\{\text{Spat-Attn}(Q[i]&,(F^{static}_{m}[i],F^{wrist}_{m}[i]))\}_{i=0}^{T}\\ Q^{\prime\prime}=\text{FFN}&(\text{Temp-Attn}(Q^{\prime})).\end{split}$ (6)
In real-world dexterous hand manipulation tasks, where $a=\{a^{xyz}\in R^{3},a^{rot}\in R^{3},a^{finger}\in R^{12}\}$,we use coefficients to balance the loss contributions from end-effector movement, rotational actions, and finger movements. Therefore, the optimization loss function for the diffusion policy can be written as:
$\begin{split}\mathcal{L}_{\text{policy}}(\psi;A)=&\omega_{xyz}\mathcal{L}_{\text{diff}}(\psi;a^{xyz})+\omega_{rot}\mathcal{L}_{\text{diff}}(\psi;a^{rot})\\ +&\omega_{finger}\mathcal{L}_{\text{diff}}(\psi;a^{finger})\end{split}$ (8)
Category Method Annotated Data Tasks completed in a row
@@
Direct Action Learning Method RT-1 [7] 100%ABC 0.533 0.222 0.094 0.038 0.013 0.90
Diffusion Policy [16] 100%ABC 0.402 0.123 0.026 0.008 0.00 0.56
Robo-Flamingo [32] 100%ABC 0.824 0.619 0.466 0.331 0.235 2.47
Future Prediction Related Method Uni-Pi [18] 100%ABC 0.560 0.160 0.080 0.080 0.040 0.92
MDT [47] 100%ABC 0.631 0.429 0.247 0.151 0.091 1.55
Susie [5] 100%ABC 0.870 0.690 0.490 0.380 0.260 2.69
GR-1 [51] 100%ABC 0.854 0.712 0.596 0.497 0.401 3.06
3D Method 3D Diffuser Actor [30] 100%ABC 0.938 0.803 0.662 0.533 0.412 3.35
Ours VPP (ours) 100%ABC 0.957 0.912 0.863 0.810 0.750 4.29
Data Efficiency MDT [47] 10%ABC 0.408 0.131 0.034 0.008 0.001 0.58
GR-1 [51] 10%ABC 0.672 0.371 0.198 0.108 0.069 1.41
@@
Table 2: Multi-task performance on Metworld. We use a single language-conditioned policy to solve all 50 tasks.

### v2

@@ Abstract
Visual representations play a crucial role in developing generalist robotic policies. Previous vision encoders, typically pre-trained with single-image reconstruction or two-image contrastive learning, tend to capture static information, often neglecting the dynamic aspects vital for embodied tasks. Recently, video diffusion models (VDMs) demonstrate the ability to predict future frames and showcase a strong understanding of physical world. We hypothesize that VDMs inherently produce visual representations that encompass both current static information and predicted future dynamics, thereby providing valuable guidance for robot action learning. Based on this hypothesis, we propose the Video Prediction Policy (VPP), which learns implicit inverse dynamics model conditioned on predicted future representations inside VDMs. To predict more precise future, we fine-tune pre-trained video foundation model on robot datasets along with internet human manipulation data. In experiments, VPP achieves a 18.6% relative improvement on the Calvin ABC-D generalization benchmark compared to the previous state-of-the-art, and demonstrates a 31.6% increase in success rates for complex real-world dexterous manipulation tasks. For your convenience, videos can be found at https://video-prediction-policy.github.io.
Machine Learning, ICML
Figure 1: Visual representations inside video prediction models explicitly express both current and future frames, providing valuable future information for embodied agent. Previous vision encoders did not have explicit future representations.
@@ 1 Introduction
Building generalist robot policies capable of solving a variety of tasks is a rapidly advancing area of research (Brohan et al., 2023; Team et al., 2024; Wu et al., 2023a; Guo et al., 2025). A crucial component in these generalist policies is the vision encoder, which captures visual information from pixel observations. Many studies have focused on optimizing vision representations for embodied agents, often leveraging internet video datasets (Ebert et al., 2021; Grauman et al., 2022) and self-supervised techniques such as single-image reconstruction (Majumdar et al., 2023; Karamcheti et al., 2023; Gupta et al., 2024), two-image contrastive learning, and image-text contrastive learning (Nair et al., 2022; Ma et al., 2022). Although these visual pre-training methods have demonstrated success for embodied tasks, they may not fully exploit the dynamic information encoded in sequential video datasets, as they typically operate on only one or two sampled images.
Recently, powerful video diffusion models (VDMs) (Ho et al., 2022; Blattmann et al., 2023a; Hong et al., 2022; Yang et al., 2024; Brooks et al., 2024) have achieved impressive results in video generation tasks. Instead of performing pre-training operation on single image or pairs of images, VDMs directly model entire video sequences. Text-guided video prediction models (TVPs) (Gu et al., 2023; Chen et al., 2023) can even predict future frames based on current observations and instructions, demonstrating a good understanding of the physical dynamics.
@@
Inspired by the strong prediction capabilities of TVP models, we hypothesize that they inherently contain valuable physical dynamics knowledge and can produce more effective visual representations for embodied agent. We take a deeper look at the visual representation inside TVP models. These representations are typically structured as a tensor with dimensions $(T,H,W)$, explicitly representing $1$ current step and $(T-1)$ predicted future steps (Blattmann et al., 2023a), where $H$ and $W$ correspond to the height and width of the image representation. In contrast, previous vision encoders do not explicitly capture future representations, as shown in Figure 1. Based on this distinction, we refer to these latent variables within the video diffusion model as “predictive visual representations”.
Our key insight is that the downstream policy can implicitly learn the inverse dynamics model by tracking the robot’s movements within the predictive representation. As long as the video model accurately predicts future scenarios for diverse tasks, the policy can generate appropriate actions by tracking robot arm’s position implicitly. In this way, we can transfer the generalization capabilities of the video prediction model to robotic policy. We only need few demonstrations to align the robot’s action space with the visual space.
Building on this insight, we introduce the Video Prediction Policy (VPP), which employs a two-stage learning process: First, we fine-tune a general-purpose video diffusion model into a text-guided video prediction (TVP) model using internet human and robot manipulation data (Goyal et al., 2017; O’Neill et al., 2023). This step aims to develop a controllable video generation model that improves prediction capabilities in the manipulation domain. In the second stage, we learn a inverse dynamics model conditioned on the predictive representations from the TVP model. Since we direct use the internal representation and avoid the need for multiple denoising steps as in previous work (Black et al., 2023; Du et al., 2024), VPP can operate at high frequency in a closed-loop manner. We also visualize the representations within the VDM and confirm that they effectively capture key information about future evolution.
@@ 2 Related Works
Visual Representation Learning for Robotics. Self-supervised learning (SSL) techniques, such as contrastive (Chen et al., 2021, 2020), distillation-based  (Baevski et al., 2022; Caron et al., 2020), and reconstructive (He et al., 2022; Bao et al., 2021), have achieved significant advancements in visual representation learning. Prior research has shown that these SSL techniques enable vision encoders to produce effective representations for embodied AI tasks (Yadav et al., 2023b, a; Parisi et al., 2022; Radosavovic et al., 2023; Chen et al., 2024a), capturing both high-level semantic and low-level spatial information. Notably, methods like R3M (Nair et al., 2022), vip (Ma et al., 2022), VC-1 (Majumdar et al., 2023), and Voltron (Karamcheti et al., 2023) have specifically focused on embodied tasks by innovating pre-training approaches on human manipulation video datasets (Goyal et al., 2017; Grauman et al., 2022). However, regardless of the training objective, the learned vision encoders primarily focus on extracting pertinent information from current observations without explicitly predicting future states. In contrast, our Video Prediction Policy leverages predictive representations within video prediction models to explicitly encapsulate both current and predicted future frames.
@@
Future Prediction for Embodied Control Tasks. Existing research also explores the use of future prediction to enhance policy learning (Bharadhwaj et al., 2024; Chen et al., 2024b; Ye et al., 2024; Guo et al., 2024). For example, SuSIE (Black et al., 2023) conditions its control policy on a predicted future keyframe generated by InstructPix2Pix (Brooks et al., 2023), while UniPi (Du et al., 2024) learns the inverse dynamics between two generated frames. These methods rely on a single future prediction step to determine actions, which may not accurately capture the complexities of physical dynamics. Additionally, they denoise the final future image which is time-cosuming and lead to low control frequency. GR-1 (Wu et al., 2023a) generates subsequent frames and actions autoregressively. However, it only generates one image per forward pass, and its prediction quality lags behind that of diffusion-based methods. Furthermore, GR-1 does not leverage pre-trained video foundation models. In contrast, VPP leverages representation fine-tuned from video foundation model, and predict a sequence of future frames to more effectively inform policy learning.
@@
Visual Representation inside Diffusion Models. Diffusion models have achieved remarkable success in the image and video generation tasks (Rombach et al., 2022; Blattmann et al., 2023a). Although diffusion models are trained as denoisers, researches have shown that image diffusion models can also function effectively as vision encoders, generating meaningful visual representations that is linear-separable for discrimination tasks (Xiang et al., 2023) and invaluable for semantic segmentation (Luo et al., 2024). Gupta et al. (2024) also point out that representation inside image diffusion are versatile for embodied tasks. However, the capabilities of representations within video diffusion models have not been extensively explored. Our findings suggest that representation within VDMs have a unique predictive property, making them especially useful for sequential embodied control tasks.
@@ 3 Preliminaries
Diffusion Policy. The diffusion model has also proven effective in action learning, known as diffusion policy (Chi et al., 2023). The diffusion policy aims to denoise the action sequence $a_{i}=(\widehat{a}_{i},\widehat{a}_{i+1},...,\widehat{a}_{i+m})$ based on observations $s_{i}$ and instruction. Chi et al. (Chi et al., 2023) point out that diffusion policy is capable of expressing complex multimodal action distributions and stabilizing training. Recent work (Reuss et al., 2024) further enhances the diffusion policy by incorporating the advanced diffusion transformer (DiT) block (Peebles & Xie, 2023), a technique we also adopt in the Video Prediction Policy to improve performance.
Figure 2: In the first stage, VPP fine-tunes a general-purpose video foundation model into a manipulation-focused Text-guided Video Prediction (TVP) model using robot and internet manipulation datasets. In the second stage, we use video-former to aggregate the representations from the TVP model during the first forward pass, followed by the diffusion policy head. This approach enables VPP to learn an implicit inverse dynamics model from the predicted future while maintaining a high control frequency.
@@ 4 Video Prediction Policy
We chose the open-sourced Stable Video Diffusion (SVD) model (Blattmann et al., 2023a) with 1.5 billion parameters as our foundation. we observe that the open-sourced SVD model conditions only on initial-frame images $s_{0}$. We augment the model to incorporate CLIP (Radford et al., 2021) language feature $l_{emb}$ using cross-attention layers. Furthermore, we adjust the output video resolution to 16×256×256 to improve training and inference efficiency. Despite these modifications, we preserve the other components of the original pre-trained SVD framework to retain its core capabilities. We denote this modified version as $V_{\theta}$. In this setup, the initial observation $s_{0}$ is concatenated channel-wise with each predicted frame as a condition. Then model $V_{\theta}$ is trained with diffusion objective, reconstructing the full video sequence $x_{0}=s_{0:T}$ in dataset $D$ from noised samples $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$:
@@
The video prediction objective offers a unified interface that directly generates future visual sequences, enabling the TVP model to harness physical knowledge from diverse datasets. These include internet human manipulation datasets $D_{H}$, internet robot manipulation data $D_{R}$, and also self-collected datasets $D_{C}$. Given the varying quality and scale of these datasets, we introduce specific coefficients $\lambda$ to appropriately balance the influence of different dataset types:
@@ 4.2 Action Learning Conditioned on Predictive Visual Representation
TVP Model as Vision Encoder. After training the TVP model specifically for manipulation tasks, it can accurately predict future sequences based on image observations and instructions. However, denoising an entire video sequence is highly time-consuming and may lead to open-loop control issues, as discussed in (Du et al., 2024). Moreover, videos in their original pixel format often contain excessive, irrelevant information that can interfere with effective decision-making.
To address these concerns, we employ the video diffusion model primarily as a “vision encoder” rather than a “denoiser” by performing only a single forward step. Our insight is that the first forward step, while not yielding a clear video, still provides a rough trajectory of future states and valuable guidance. This insight is verified in our experiment section and shown in Fig 4. Specifically, we concatenate the current image $s_{0}$ with the final noised latent $q(x_{t^{\prime}}|x_{0})$ (typically white noise) and input this combination into the TVP model. We then directly leverage the latent features. Previous work (Xiang et al., 2023) highlights that up-sampling layers in diffusion models yield more effective representations. The feature at the $m^{th}$ up-sampling layer, with width $W_{m}$ and height $H_{m}$, is expressed as:
$L_{m}=V_{\theta}(x_{t^{\prime}},l_{emb},s_{0})_{(m)},L_{m}\in\mathbb{R}^{T\times C_{m}\times W_{m}\times H_{m}}$
To effectively aggregate features from the up-sampling layers and eliminate the need for manual layer selection, we propose an automatic method for aggregating features across different layers. First, we linearly interpolate each layer’s feature map to the same height and width $W_{p}\times H_{p}$:
$L_{m}^{\prime}=\text{Interpolation}(L_{m}),L_{m}^{\prime}\in\mathbb{R}^{T\times C_{m}\times W_{p}\times H_{p}}$
We then stack the features along the channel dimension. The final predictive visual representation $F_{p}\in\mathbb{R}^{T\times(\sum_{m}C_{m})\times W_{p}\times H_{p}}$ is given by:
$F_{p}=\text{concate}((L_{0}^{\prime},L_{1}^{\prime},\dots,L_{m}^{\prime}),dim=1)$
For a robot with multiple camera views, such as a third-view and a wristed camera, we predict the future for each view independently, denoted as $F^{static}_{p},F^{wrist}_{p}$.
Video Former. These predictive representations within the video diffusion model are still high-dimensional, as they express a sequence of image features. To efficiently aggregate representations across spatial, temporal, and multi-view dimensions, we design a Video Former to consolidate this information into a fixed number of tokens. The Video Former initializes learnable tokens $Q_{[0:T,0:L]}$ with fixed length $T\times L$, performing spatial-temporal attention (Blattmann et al., 2023b) on each corresponding frame, followed by feed-forward layers. Formally, this branch can be expressed as follows where $i$ is the index of frame:
$\begin{split}Q^{\prime}=\{\text{Spat-Attn}(Q[i]&,(F^{static}_{p}[i],F^{wrist}_{p}[i]))\}_{i=0}^{T}\\ Q^{\prime\prime}=\text{FFN}&(\text{Temp-Attn}(Q^{\prime})).\end{split}$ (5)
Category Method Annotated Data $i^{th}$ Task Success Rate
@@
Direct Action Learning Method RT-1 100%ABC 0.533 0.222 0.094 0.038 0.013 0.90
Diffusion Policy 100%ABC 0.402 0.123 0.026 0.008 0.00 0.56
Robo-Flamingo 100%ABC 0.824 0.619 0.466 0.331 0.235 2.47
Future Prediction Related Method Uni-Pi 100%ABC 0.560 0.160 0.080 0.080 0.040 0.92
MDT 100%ABC 0.631 0.429 0.247 0.151 0.091 1.55
Susie 100%ABC 0.870 0.690 0.490 0.380 0.260 2.69
GR-1 100%ABC 0.854 0.712 0.596 0.497 0.401 3.06
Vidman 100%ABC 0.915 0.764 0.682 0.592 0.467 3.42
3D Method RoboUniview 100%ABC 0.942 0.842 0.734 0.622 0.507 3.65
Ours VPP (ours) 100%ABC 0.965 0.909 0.866 0.820 0.769 4.33
Data Efficiency GR-1 10%ABC 0.672 0.371 0.198 0.108 0.069 1.41
@@
Table 2: Multi-task success rate on Metaworld. We use a single language-conditioned policy to solve all 50 tasks.